Oscar Clemente is the manager of Forbes Division of Pitt, Inc., a manufacturer of biotech products. Forbes Division, which has $3 million in assets, manufactures a special testing device. At the beginning of the current year, Forbes invested $3 million in automated equipment for test machine assembly. The division's expected income statement at the beginning of the year was as follows.

 





Sales revenue$15,500,000
Operating costs


Variable
2,050,000
Fixed (all cash)
7,500,000
Depreciation


New equipment
1,480,000
Other
1,200,000
Division operating profit$3,270,000

 

A sales representative from LSI Machine Company approached Oscar in October. LSI has for $5.75 million a new assembly machine that offers significant improvements over the equipment Oscar bought at the beginning of the year. The new equipment would expand division output by 10 percent while reducing cash fixed costs by 5 percent. It would be depreciated for accounting purposes over a 3-year life. Depreciation would be net of the $500,000 salvage value of the new machine. The new equipment meets Pitt's 20 percent cost of capital criterion. If Oscar purchases the new machine, it must be installed prior to the end of the year. For practical purposes, though, Oscar can ignore depreciation on the new machine because it will not go into operation until the start of the next year.

 

The old machine, which has no salvage value, must be disposed of to make room for the new machine.

 

Pitt has a performance evaluation and bonus plan based on ROI. The return includes any losses on disposal of equipment. Investment is computed based on the end-of-year balance of assets, net book value. Ignore taxes.

 

Required:

a. What is Forbes Division's ROI if Oscar does not acquire the new machine? (Enter your answer as a percentage rounded to 1 decimal place (i.e., 32.1).)

b. What is Forbes Division's ROI this year if Oscar acquires the new machine? (Enter your answer as a percentage rounded to 1 decimal place (i.e., 32.1).)

c. If Oscar acquires the new machine and it operates according to specifications, what ROI is expected for next year? (Enter your answer as a percentage rounded to 1 decimal place (i.e., 32.1).)
Answer #1

A) 3,270,000/ (3,000,000+ 3,000,000- 1,480,000- 1,200,000)= 0.984939759 or

Answer: 98.5% 


B) step 1) 3,000,000- 1,480,000= 1,520,000

step 2) 3,270,000- 1,520,000= 1,750,000

3) 3,000,000- 1,200,000+ 5,750,000= 7,550,000

4) 1,750,000/ 7,550,000= 0.2317880795 or Answer: 23.2 %


C) revenue  15,500,000* 10%= 1,550,000+ 15,500,000= 17,050,000

variable  2,050,000* 10%= 205,000+ 2,050,000= 2,255,000

fixed  7,500,000* 5%= 375,000- 7,500,000= 7,125,000

equipment:  (5,750,000(from 5.75) -500,000)/3 year-life= 1,750,000

other 1,200,000 same as in question

operating profit 17,050,000- 2,255,000- 7,125,000- 1,750,000- 1,200,000= 

4,720,000


3,000,000- (2* 1,200,000)+ 5,750,000- 1,750,000= 4,600,000


4,720,000/ 4,600,000= 1.0260869565 or Answer: 102.6 %